Download
general status of commissioning n.
Skip this Video
Loading SlideShow in 5 Seconds..
General Status of Commissioning PowerPoint Presentation
Download Presentation
General Status of Commissioning

General Status of Commissioning

128 Vues Download Presentation
Télécharger la présentation

General Status of Commissioning

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. General Status ofCommissioning Piotr Krzysztof Skowroński CTF3 Collaboration Technical Meeting 27January 2009

  2. Commissioning Proving that machine is technically operational and the beam is as we want it to be • Outline • Status year ago • Operations and machine setup • Activities in 2008 • Current status on optics control 2

  3. recombinationx 4 recombination x 2 bunch lengthcontrol bunchcompression fully loadedacceleration structures12 GHz phase-coding PETS Decelerationstability two-beamacceleration structures 30 GHz CTF3 base line programme CTF3 – R&D Issues - where

  4. CRM TL1 DL CR TL2 CLEX CLIC Test Facility 3 TL1 & CRM commissioned fall 2006 Delay Loop: commissioned with beam 2005-2006 Linac: commissioned with beam 2003 - 2004 CR commissioningsince 2007 TL2 and TBTS commissioning started August 2008 4

  5. Main Points of 2006 • Delay Loop Commissioned • 7A after DL • 1400 ns pulse length • TL1 and CR injectionCommissioned • 3.2A in TL1, 3A in CRM • RF injection to CR 2 1 3 CTF3 Collaboration Meeting

  6. Main Points of 2007TL1 and CR • Transfer Line 1 relatively quickly commissioned • No major problems with establishing injection to CR • Long and painful CR commissioning being finished only now • Fast Vertical Instability caused by the deflectors • Bad cabling of quadrupoles • Wrongly calibrated BPMs • Unstable gun • Inaccurate model of the ring current Vpos Hpos 6

  7. Main Points of 2007CR results • Besides all the problems • We have managed to recombine the train factor 4 • OK, almost 4, but the principle was shown to be valid • Perform the optics studies which enabled gradualdebugging and validation of the machine model CTF3 Collaboration Meeting

  8. Operations • The machine is operated in two modes • PETS beam for 30GHz structures tests • From Friday to Monday morning • Setup is done by CTF3 team • Night and weekend supervision from CCC by PS operators • Commissioning beam • From Monday to Thursday • Two overlapping shifts (voluntary), • From 8AM to 4PM • From 2PM to 8PM

  9. Machine Setup on one slide Machine setup is based on quad scans Beam is set-up to the screen on girder 10 Phases of acc-structures up to girder 10 are adjusted to get full loading Twiss parameters are measured and whole linac is rematched Phase adjustment for the rest of the structures Energy profile is measured at girder 10 and elsewhere is calculated Another quad scan is made at CTS TL1 and is matched to the CR closed solution at this given energy If something changes at given point, we must redo the setup from there CRM TL1 DL CR TL2 CLEX

  10. Beam recombination in CR is made with time variable bump injection line septum 1st deflector 2nd deflector local inner orbits RF deflector field lo CR Injection (b) (a) 10 CTF3 Collaboration Meeting 27 Jan 2009

  11. Long circulating CR beam Many measurements require long circulating beam in CR Tune Closed orbit correction Closed orbit length measurement The beam is injected to the ring using RF deflector The klystron timing is adjusted so the RF pulse ends when the last bunch of the train leaves the deflector Since those are travelling wave cavities with short filling time the RF is gone before the first beam of the train comes back to the deflector This way beam makes several hundreds turns before it irradiates too much energy thru synchrotron radiation to be fitted in the acceptance

  12. 2008Schedule 30 GHz DL & CR 30GHz only RUN 1 DL, CR, TL2 TBTS Installation:TL2, CALIFES, TBTS, new deflectors Today Linac only RUN 2 Linac, Ring area, (CLEX) PETS in TBTS CALIFES Installation: tail clipper RUN 3 CTF3 Collaboration Meeting

  13. 2008 program – Run 1 • Find and repair the problem with the gun • Linac • Establish stable & documented working point • Stability studies • diagnostic consolidation • Delay Loop • Beam optics measurements • Re-establish recombination • Finish the commissioning of CR • Perform set of optics measurements • Validate the machine model • Tune and bfunction dependence of vertical instability • Test new devices and techniques • EuroTeV BPMs • Bunch length measurement with RF pickup • Transient compensation • Dispersion Free Steering and other automatic algorithms CTF3 Collaboration Meeting

  14. Unstable gun Unstable Stable 14 CTF3 Collaboration Meeting • Already since the first run of 2007 we observe large beam jitter that originates from the gun instability • The current is slightly different from shot to shot • Since we operate in the fully loaded mode • phases are adjusted so all the RF energy is transferred to the beam bunches with different current are accelerated differently • It leads to position jitter in the dispersive regions • What makes the setup and measurements very hard! 27 Jan 2009

  15. Unstable gun (cont.) 15 CTF3 Collaboration Meeting • Finding the problem was difficult and tedious • The system operates on high voltage (130-150kV) and hence specialized knowledge and equipment is needed for investigations • On top of that the device is constructed by LAL and the designer is retired • It was finally found that the high voltage power converter produces chainsaw-like signal • however according to the manufacturer the device itself was OK • After inspection in-situ one of resistors was found cracked • connecting capacitor bank to the ground • at low voltage the resistance was 0, but when 150kV applied than it was conducting thru arcing • And later another resistor found cracked • between capacitor bank and the high voltage deck • Since than (end of September 2008) the gun behaves OK • However, we are not certain that it was the core of the problem and it is going to come back at some point 27 Jan 2009

  16. Difficult Run 1 • Due to the gun problems, but also many other technical problems like • Instable water temperatures for the pulse compression cavities • Controls • Timing • VME crates overloaded and hanging up permanently • Vacuum leaks • Klystrons often down the first run was not extremely productive • Anyway me managed to • Perform optics measurements and study Linac and TL1 • Response Matrix • Dispersion • Consolidate beam diagnostics in those sections

  17. Bunch length measurements CTF3 Collaboration Meeting • We have two methods to measure the bunch length • Standard way with RF deflector and MTV screen • Novel non-destructive RF pickup • For further details see Anna Dabrowskitalk at 3PM

  18. Steering Algorithms Linac Orbit after all-to-all correction Comparison between results of Dispersion Free Steering and All-to-All CTF3 Collaboration Meeting • We have successfully, but not without problems applied following steering techniques • dispersion free • one-to-one • all-to-all

  19. EuroTeV BPMs Stdev1PBPM _1.5A= 0.6µm 19 CTF3 Collaboration Meeting • We have tested the new high precision BPMs • We have already tried during previous year however, we could not get full transmission thru 2nd and 3rd BPM • Turned out to be a hardware problem • We have installed standard BPM after the tested triplet • Measurements successfully repeated this year 27 Jan 2009

  20. CRM TL1 DL CR TL2 CLEX 2008 program – Installations 1 • Technical break in June and July for • Complete TL2 installation • Install new deflectors • Complete CLEX installations • When the dump after CR was removed to make place for TL2 personnel was not allowed to enter CLEX when the beam was in DL or CR (and vice versa) CTF3 Collaboration Meeting

  21. CRM TL1 DL CR TL2 CLEX 2008 initial program for Run 2 • Run 2 • Commission the new RF deflectors • Re-setup beam recombination in CR • Full recombination with DL and CR (factor 8) • Commissioning of • TL2 • TBTS (no PETS) • First trial for CALIFES • Water leak on a bi-metal joint • Coherent Diffraction Radiation measurements CTF3 Collaboration Meeting

  22. TL2 and TBTS Commissioning Current Horizontal Vertical 22 CTF3 Collaboration Meeting • Strip-line kicker works well • BPMs based on the new LAPP electronics in principle work, but • 30ns time jitter • Big problems with OASIS viewer • The traces are digitized in-situ and analog signals are not available and we have to use this application to get them • Swapped vertical with horizontal for one type of pickups, OK for the other • Front-end task often crashing • Constant improvement in making them stably operational • See F.Tecker talk at 14.40 for more • No MTV suitable for quad scans • The beam setup was very,very, very difficult 27 Jan 2009

  23. TL2 and TBTS Commissioning 23 CTF3 Collaboration Meeting • After long fight we have managed to transport the beam till the end of TBTS, however, with 50% current loss • We hadsome problem with the optics, which was not diagnosed yet • Either incoming CR beam wasterribly mismatched • Rather doubtful since the beam was able to circulate for tens of turns in CR • There wassome optics error in TL2 itself, but it was difficult to diagnose not having • Full transmission • Operational beam monitors 27 Jan 2009

  24. TL2 and TBTS Commissioning We have attempted to do some optics measurements, but they were not very successful mainly due to Losses along the line One has to get somehow full transmission and only then try optics measurements Problems with BPMs

  25. New deflectors in Combiner Ring The Instability Gone current vertical horizontal Old deflectors, undamped BPM signals shoving vertical instability New deflectors, damped damping loads 25 CTF3 Collaboration Meeting 27 Jan 2009 BPM signals, no instability

  26. New deflectors in Combiner Ring The Instability Gone 26 CTF3 Collaboration Meeting • Stable beam circulating for miliseconds • Recombination factor 4 • Since than all the measurementswere much easier to do andmuch more precise • Tune measurements • Orbit corrections • Still the model did notdescribe the ring opticsprecisely enough 27 Jan 2009

  27. Bunch length from Coherent Diffraction Radiation Prototype device installed in CRM line Initially very noisy signals Improved device is much better Measured radiation in function of target position consistent with the expectations Good start for this year See PavelKarataev talk today at 5PM

  28. CRM TL1 DL CR TL2 CLEX 2008 Program • Installations in October • Tail clipper • CALIFES installations • PETS in TBTS • Initial program for Run 3 • Commissioning of • CALIFES • TBTS (with PETS) • PHIN CTF3 Collaboration Meeting

  29. RF power from PETS in TBTS Power produced by PETS (Ibeam=2A)without and with recirculation

  30. Recombined beam to PETS • 8A through PETS • Almost 30MW obtained

  31. PHIN Commissionning First measurements in Nov Nominal 1.3 ms pulse with 666 ps bunch separation Measured nominal energy with good energy spread 5.3 MeV with 0.5 % σ energy spread Emittance ‘to good to be true’ Laser control and diagnostic insufficient Strange asymmetric beam See Wednesday afternoon presentations

  32. CALIFES Commissioning Started on Dec 1 The beam with laser pulse train of 100 ns length (150 bunches) 0.073 nC per bunch 100% transmission thru the 1st cavity Problems with RF phase adjustment Finished on Dec 11 Laser broken See Wilfrid Farabolini talk on Wednesday

  33. Optics Studies andModel Validation • The main reason for extensive optics measurements is verification that the model predicted optics well agrees with the measured one, and if they don’t • Machine was not build according to the specification • There is a bug in the model • One can not back-hand the setup with the model calculations • We performed the following optics measurements • Dispersion • Response Matrix • Tunes • Combiner Ring Length • Bunch length • Quadrupole current to K1 coefficient CTF3 Collaboration Meeting

  34. Linac Model Validation Current Status and Remaining Problems CTF3 Collaboration Meeting

  35. Linac Model Validation Current Status and Remaining Problems CTF3 Collaboration Meeting • There is some optics error in the area of girder 7 and 8 • We see it as phase slippage in response matrix • But also when we rematch the linac, the Twiss parameters measured afterwards are slightly different from the expected ones • The magnetic fields on the pole surfaces where verified and are consistent with all the other magnets of this family • More work is needed to nail down the problem • Additional data, for example response matrix with different quad families off, would certainly help • However, in principle we have already sufficient data to be able to find it with off-line analysis and systematic analysis should reveal the problem

  36. TL1 Model Validation Discovered Bugs • The measured dispersion pattern in TL1 always was showing slight (just outside the error-bar) but systematic discrepancy at the last BPM • Finally with the high precision response matrix measurement we found • wrong current to gradient coefficient for CT.QDF0720 • Wrong drift length because of the same variable name used in CR CTF3 Collaboration Meeting

  37. TL1 Model Validation Current Status and Remaining Problems CTF3 Collaboration Meeting • We consider TL1 commissioned and its model quite well validated, although some small discrepancies remain

  38. CR Model ValidationDiscovered Bugs • Already during 2007 runs response matrix study allowed to find few problems with the CR model • Current to gradient coefficient for J type magnets (4%) • Cross-cabled control units for quadrupole’s power converters • Some discrepancies were still present • They were not localized in a precise location • Large statistical uncertainties originating in the gun instability did not allow to find the subtle problems quickly • Further higher precision measurements were needed to nail down the remaining discrepancies CTF3 Collaboration Meeting

  39. CR Model ValidationResponse Matrix Results • Since CR has two fold symmetry the responses should be symmetric • An they are (more or less)http://skowron.web.cern.ch/skowron/results/cr_respmtx/20081209CR/symm.html • There is no error due to single quad or family • Almost certainly the error has to be related to wrong magnetic coefficient of a type

  40. CR Model Validation Current Status and Remaining Problems • Lowering 10% quadrupolar strength for combined function magnets • and • Increasing focusing of J type by 5.5% • gives good agreement between the model and RM data ... • http://skowron.web.cern.ch/skowron/results/cr_respmtx/20081209CR/hor.html • http://skowron.web.cern.ch/skowron/results/cr_respmtx/20081209CR/ver.html • ... and with other measurements CTF3 Collaboration Meeting

  41. CR Model Validation Dispersion measurements CTF3 Collaboration Meeting

  42. CR Model Validation Tune measurements 42 CTF3 Collaboration Meeting 27 Jan 2009

  43. CR Model Validation Old Response Measurements • At different energies and optics 85MeV August 2008 104MeV November 2007

  44. CR Model Validation Current Status and Remaining Problems • However • Does not describe some of old (with the old deflectors) tune measurements • It is hard to believe that focusing of this magnets is not what expected • They were used in EPA • Measured before installation in EPA and before installation in CTF • And the results are coherent • Focusing in this case is an strictly geometric issue • For given bending angle it must have given k1 • We have 3 of this magnets in TL1, and over there we do not see any problems • We need to spend some more time on this problem 44 CTF3 Collaboration Meeting 27 Jan 2009

  45. Orbit corrections Before the correction After the correction CTF3 Collaboration Meeting • Successful application of orbit correction in CR • See the next talk for more details

  46. TL2 model validationMeasurements • We have attempted to do the optics measurements in TL2, however they are uncertain due to • Problems with BPM readings • Limited dynamic aperture • The optics was set „by hand” • Any „kick” or „scaling” was leading to the beam losses • Within the precision we have we do not see any optics error

  47. Conclusions • 2008 runs went relatively smoothly without any major disasters like • Vacuum leaks (2 in 2007) • Klystrons down for good (2 in 2007) • It seems the problem with the unstable gun disappeared for now • But we are not really sure we have found the core of the problem and that it disappeared for good • New RF deflectors installed in the Combiner Ring • It is confirmed now that there is no sign of the instability now • Immediately managed to get recombination factor 4 • There is serious chance to finish its commissioning soon • TL2 and TBTL need more reliable diagnostic for optics measurements • CALIFES commissioning postponed to the next run CTF3 Collaboration Meeting

  48. Conclusions • The gun • Find and repair the problem that makes unstable • Consolidation program needs to be completed • Linac • Establish stable & documented working point • Stability studies • Diagnostic consolidation • TL1 • Optics measurements and model validation • Delay Loop • Beam optics measurements • Re-establish the recombination CTF3 Collaboration Meeting

  49. Conclusions • Finish the commissioning of CR • Perform set of optics measurements • Validate the machine model • Tune and bfunction dependence of vertical instability • Problem disappeared with the new deflectors • Commissioning • TL2 • TBTS • TBL • CALIFES • Test new devices and techniques • EuroTeV BPMs • Bunch length measurement with RF pickup • Transient compensation • Dispersion Free Steering and other automatic algorithms CTF3 Collaboration Meeting

  50. 2008 structures 12 GHz PETS on-off deceleration stability two-beam acceleration structures 30 GHz 2006-2008 recombination x 4 recombination x 2 2005 bunch length control bunch compression fully loaded acceleration 2004 phase-coding 2009 2010